# A Structured Framework for Evaluating and Enhancing Interpretive Capabilities of Multimodal LLMs in Culturally Situated Tasks

### v1

@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable performance on general NLP benchmarks, yet their applicability in culturally embedded, humanistic domains remains limited. In high-context interpretive tasks such as art criticism, clinical narrative analysis, or historical commentary, model performance depends not only on linguistic fluency or factual accuracy, but also on deeper forms of cognitive alignmentepistemic sensitivity, rhetorical coherence, and cultural adaptability.
@@
Meanwhile, humanistic commentary often exhibits non-linear logic, specialized lexicons, and varied stylistic conventions, particularly in Chinese art contexts where rhetorical strategies such as yijing (yijing, artistic conception) or qiyun shengdong (qiyun shengdong, spiritual resonance) are essential but difficult to quantify Bush (1971); Sirén (1936). Without appropriate grounding, LLMs risk producing synthetic outputs that mimic surface patterns but fail to demonstrate epistemic alignment Guo et al. (2023). This growing mismatch calls for new paradigms in evaluation and adaptation.
To address these challenges, we introduce VULCAthe Vision-Understanding and Language-based Cultural Adaptability Framework. VULCA is a structured evaluation and enhancement framework designed to assess how well VLMs align with domain-specific interpretive practices in culturally situated tasks. Our work centers on Chinese art commentary, but the methodology generalizes to other multimodal and epistemically rich domains such as religion, medicine, or history. VULCA combines three core components: (1) a multi-dimensional human expert benchmark (MHEB) constructed from 163 art commentaries annotated across five cultural capability dimensions; (2) a persona-guided recontextualization mechanism using eight interpretive personas and a domain-specific knowledge base; and (3) a joint evaluation pipeline integrating vector-space semantic alignment with rubric-based capability scoring. Commentaries are generated from annotated traditional Chinese paintings, and their alignment with expert patterns is evaluated with and without interventions. As a result, we produce five contributions: (i) the definition of VULCA, a new structured framework for assessing and enhancing VLMs in culturally grounded, multimodal reasoning tasks; (ii) we construct MHEB, a high-quality human benchmark of Chinese art commentary annotated across five capability dimensions; (iii) we develop and evaluate persona-guided recontextualization interventions using eight expert personas and a domain-specific knowledge base; (iv) we demonstrate over 20% improvement in symbolic reasoning and over 30% improvement in argumentative coherence on Gemini 2.5 Pro using our proposed method; and (v) we establish the generalizability of our evaluation methodology to other epistemically rich domains such as religion, history, and education.
@@ 2 Related Work
Despite significant advances in multimodal evaluation, current benchmarks primarily target factual understanding rather than cultural interpretation. Existing benchmarks for large or multimodal language models, such as Fu et al. (2024); You et al. (2023), emphasize factual accuracy or instruction following, seldom addressing symbolic interpretation or epistemic alignment. Recent cultural evaluation efforts like M3Exam Zhang et al. (2023) and SEED-Bench Li et al. (2024) begin to incorporate cultural knowledge but focus on factual recall rather than interpretive reasoning. ArtGPT Yuan et al. (2024), for instance, evaluates stylistic generation but lacks formal metrics for interpretive depth. While prior work explores aesthetic reasoning Wang (2024), these studies rarely offer structured, multi-capability evaluation. Our work addresses this gap by introducing cultural adaptability, operationalized through a multi-dimensional human expert benchmark with capability rubrics, enabling quantitative comparison in high-context domains like Chinese art.
@@
Gap in Multimodal InputInterpretation Evaluation.
@@
Surveys Guo et al. (2023) discuss LLM limitations in nuanced discourse, but few studies compare model responsiveness to structured cultural interventions. Our empirical evaluation shows persona and knowledge base intervention improves symbolic reasoning and argumentative coherence by over 2030%, highlighting epistemic alignment’s role beyond fluency. This cross-model, capability-specific analysis distinguishes our work.
@@ 3.1 MHEB Construction and Annotation Process
(1) Chinese Art Theory: Building on Xie He’s Six Canons (liufa, 550 CE) Xie (550), particularly the concepts of “spirit resonance” (qiyun shengdong) and “bone method” (gufa yongbi), which inform our Feature Focus dimension’s emphasis on brushwork technique, artistic conception, and emotional expression.
@@
Data Collection. We collected 163 expert commentaries from authoritative sources including museum catalogs from the Palace Museum Beijing, National Palace Museum Taipei, and Metropolitan Museum of Art, as well as peer-reviewed art history journals and monographs by recognized scholars specializing in Qing court painting. Each commentary averages 500-800 Chinese characters and provides in-depth analysis of specific paintings from the “Twelve Months” series. The annotation process generated 558 total annotation instances (163 texts 3 annotators plus quality control samples), which were consolidated into 163 final records after resolving disagreements.
Expert Sources. The 163 commentaries in MHEB were extracted from scholarly publications by 9 distinguished art historians specializing in Chinese painting and Qing court art. The corpus includes: Xue Yongnian (17 texts from two monographs), Wang Di (28 texts), Yang Danxia (28 texts), Nie Chongzheng (15 texts), Shan Guoqiang (18 texts), Li Shi (17 texts), Xu Jianrong (17 texts), Zhu Wanzhang (11 texts), and Chen Yunru (12 texts). These experts represent major institutions including the Palace Museum Beijing, National Palace Museum Taipei, and leading Chinese art history departments, ensuring diverse yet authoritative perspectives on Giuseppe Castiglione’s “Twelve Months” series.
@@
Inter-Annotator Agreement (IAA). To quantitatively assess the reliability of our annotations, we calculated inter-annotator agreement using two complementary metrics. For categorical labels (e.g., stance categories), we computed Fleiss’ kappa Fleiss (1971), which measures agreement beyond chance for multiple raters. For continuous scores (e.g., feature prominence ratings from 0-1), we calculated the intraclass correlation coefficient (ICC) Shrout and Fleiss (1979), which assesses the consistency of quantitative measurements across raters. The average Fleiss’ kappa across stance categories was 0.78, indicating substantial agreement according to Landis and Koch’s interpretation scale. The ICC for feature prominence scores reached 0.82, demonstrating excellent reliability. When disagreements occurred (defined as < 0.6 for specific labels), they were resolved through discussion, with a senior art historian serving as arbiter for persistent conflicts. The stable inter-rater agreement ( variation < 0.05 across time) validated the effectiveness of our quality control measures. Final dataset statistics show balanced representation across different evaluative stances (Historical: 31%, Aesthetic: 28%, Technical: 23%, Comparative: 18%) and comprehensive coverage of feature focus.
@@ 3.2 Feature Engineering from Human Expert Critiques
Framework foundation relies on human expert commentaries, significantly from Giuseppe Castiglione’s (Lang Shining) “Twelve Months” () seriesQing imperial court paintings fusing Chinese and Western traditions. To enhance model training and evaluation, a sliding window cropping strategy (640640 pixel sub-images) was applied to these high-resolution images, augmenting data diversity and granularity for improved VLM detail recognition and evaluation accuracy, a common practice in computer vision (e.g., Lin et al. (2014); Krishna et al. (2017)).
@@ 3.3 Evaluation of Dimensions and Label System
The three dimensions of our framework (i.e., Evaluative Stance, Feature Focus, and Commentary Quality) were derived from multiple sources: (1) traditional Chinese painting theory, particularly Xie He’s “Six Principles of Painting” () Acker (1954) which emphasizes spirit resonance (qiyun shengdong), bone method (gufa yongbi), and correspondence to nature (); (2) Western art criticism frameworks from Panofsky’s three levels of meaning Panofsky (1955) and Wlfflin’s formal analysis principles Wölfflin (1950); (3) consultations with curators from the Palace Museum and Metropolitan Museum who validated the relevance of these dimensions for Qing court painting analysis; and (4) empirical analysis of recurring patterns in our collected expert commentaries.
Evaluative Stance characterizes the rhetorical or evaluative position taken by the commentator (e.g., historical interpretation, praise, or critique). Feature Focus identifies the specific visual or contextual aspects discussed in the commentary (e.g., line quality, symbolism, spatial composition). Commentary Quality captures the analytical depth and logical structure of the commentary, ranging from clear, well-argued insights to superficial or biased remarks. Furthermore, each dimension comprises a set of fine-grained subcategories with bilingual EnglishChinese mappings. Full definitions and label lists are provided in Appendix C.5.
@@ 3.4 Construction and Definition of Critic Personas
To capture holistic critique style and depth beyond granular features, we constructed “critic personas” representing archetypal critical perspectives. Their development was data-driven, analyzing features from human expert commentaries, complemented by art history domain expertise. Five core personas were defined: Comprehensive Analyst (), Historically Focused Critic (), Technique & Style Focused Critic (), Theory & Comparison Focused Critic (), and General Descriptive Profile (). These five core personas represent data-driven evaluation categories derived from clustering analysis of human expert features, serving as benchmarks for assessing whether VLM outputs align with recognizable expert critique patterns.
@@ 3.5 Task Definition
Inputs were multifaceted: high-definition “Monthly Images” (sometimes segmented); predefined “Persona Cards” Jiang et al. (2024) serving as experimental interventions-distinct from the five evaluaction personas above, these eight cultural perspective prompts VLM generation: guiding analysisMama Zola (), Professor Elena Petrova (), Okakura Kakuz (), Brother Thomas (), John Ruskin (), Su Shi (), Guo Xi (), and Dr. Aris Thorne (); standardized prompt templates Nayak et al. (2024); and an optional JSON knowledge base Zhang et al. (2024b); Bin et al. (2024). Persona guidance aimed to assess VLM capability to simulate diverse perspectives and analytical styles Zhang et al. (2024a). See Appendix B for a detailed summary of each critic persona included in our study. To avoid confusion, we distinguish between the use of personas at two different levels: the five core personas described in the previous sub-section are data-driven evaluation categories for classifying generated critiques based on feature patterns, while the eight persona cards are cultural perspective prompts used to guide VLM generation during experiments. The former evaluates outputs, while the latter shapes inputs.
@@ 4.3 Effectiveness of Persona-Guided Interventions
1 Qwen-2.5-VL-7B + Mama Zola () + KB 9.2/10 100%
2 meta-llama_Llama-4-Scout-17B-16E-Instruct + John Ruskin () + KB 8.9/10 97%
3 meta-llama_Llama-4-Scout-17B-16E-Instruct + Mama Zola () + KB 8.7/10 95%
4 meta-llama_Llama-4-Scout-17B-16E-Instruct + Brother Thomas () + KB 8.5/10 92%
5 meta-llama_Llama-4-Scout-17B-16E-Instruct + Su Shi () + KB 8.5/10 92%
@@ A.1 Lang Shining’s "Twelve Months" Dataset
Our study centers on Giuseppe Castiglione’s “Twelve Months” series (Twelve Months), 12 paintings showing seasonal activities in the Qing imperial court. These paintings fuse Chinese and Western artistic traditions, ideal for cross-cultural interpretation study. We compiled digital images (6 million pixels) from the National Palace Museum (Taiwan) digital archives under CC BY 4.0 license. The dataset includes historical texts and scholarly analyses in both Chinese and English, from Qing Dynasty sources and modern scholarship.
@@ Appendix B Persona Definitions
B.1 Mama Zola (Mama Zola)
@@
B.2 Okakura Kakuz (Okakura Kakuzo)
@@
B.3 Professor Elena Petrova (Professor Elena Petrova)
@@
B.4 Brother Thomas (Brother Thomas)
@@
B.5 John Ruskin (John Ruskin)
@@
    • “To see clearly is poetry, prophecy, and religion, all in one.”
B.6 Su Shi (Su Shi)
@@
  • Key Influences/Background: Deeply influenced by Confucianism, Daoism, and Chan (Zen) Buddhism. Advocated for "scholar-official painting" (scholar-official painting), emphasizing the integration of poetry, calligraphy, and painting, and the expression of inner spirit. His artistic ideas had a profound impact on the development of later literati painting.
  • Analytical Style and Characteristics: Values the "spiritual resonance" (spiritual resonance) and "artistic interest" (artistic interest) of artworks over external formal likeness. Emphasizes the decisive role of the artist’s personal character, knowledge, and cultivation in creation. Esteems an aesthetic realm of natural innocence, plainness, and distanced simplicity.
@@
B.7 Guo Xi (Guo Xi)
  • Basic Information: Outstanding Chinese Northern Song Dynasty landscape painter and painting theorist (male, c. 1023-c. 1085, Wen County, Heyang). Served as an Erudite (Art Academy Scholar) in the imperial painting academy during Emperor Shenzong’s reign.
  • Key Influences/Background: Inherited and developed the traditions of the Northern school of landscape painting, emphasizing observation and experience of nature. His theoretical work "The Lofty Message of Forests and Streams" (The Lofty Message of Forests and Streams) is a seminal text in Chinese landscape painting theory.
  • Analytical Style and Characteristics: Emphasized that landscape paintings should be "walkable, viewable, wanderable, and habitable" (walkable, viewable, wanderable, habitable). Proposed methods for observing and depicting landscapes such as the "Three Distances" (Three Distances: high distance, deep distance, level distance). Valued the influence of seasons and climate on scenery, striving for majestic and varied artistic conceptions (yijing).

### v2

@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable performance on general NLP benchmarks, yet their applicability in culturally embedded, humanistic domains remains limited. In high-context interpretive tasks such as art criticism, clinical narrative analysis, or historical commentary, model performance depends not only on linguistic fluency or factual accuracy, but also on deeper forms of cognitive alignment—epistemic sensitivity, rhetorical coherence, and cultural adaptability.
@@
Meanwhile, humanistic commentary often exhibits non-linear logic, specialized lexicons, and varied stylistic conventions, particularly in Chinese art contexts where rhetorical strategies such as yijing (意境, artistic conception) or qiyun shengdong (气韵生动, spiritual resonance) are essential but difficult to quantify Bush (1971); Sirén (1936). Without appropriate grounding, LLMs risk producing synthetic outputs that mimic surface patterns but fail to demonstrate epistemic alignment Guo et al. (2023). This growing mismatch calls for new paradigms in evaluation and adaptation.
To address these challenges, we introduce VULCA—the Vision-Understanding and Language-based Cultural Adaptability Framework. VULCA is a structured evaluation and enhancement framework designed to assess how well VLMs align with domain-specific interpretive practices in culturally situated tasks. Our work centers on Chinese art commentary, but the methodology generalizes to other multimodal and epistemically rich domains such as religion, medicine, or history. VULCA combines three core components: (1) a multi-dimensional human expert benchmark (MHEB) constructed from 163 art commentaries annotated across five cultural capability dimensions; (2) a persona-guided recontextualization mechanism using eight interpretive personas and a domain-specific knowledge base; and (3) a joint evaluation pipeline integrating vector-space semantic alignment with rubric-based capability scoring. Commentaries are generated from annotated traditional Chinese paintings, and their alignment with expert patterns is evaluated with and without interventions. As a result, we produce five contributions: (i) the definition of VULCA, a new structured framework for assessing and enhancing VLMs in culturally grounded, multimodal reasoning tasks; (ii) we construct MHEB, a high-quality human benchmark of Chinese art commentary annotated across five capability dimensions; (iii) we develop and evaluate persona-guided recontextualization interventions using eight expert personas and a domain-specific knowledge base; (iv) we demonstrate over 20% improvement in symbolic reasoning and over 30% improvement in argumentative coherence on Gemini 2.5 Pro using our proposed method; and (v) we establish the generalizability of our evaluation methodology to other epistemically rich domains such as religion, history, and education.
@@ 2 Related Work
Despite significant advances in multimodal evaluation, current benchmarks primarily target factual understanding rather than cultural interpretation. Existing benchmarks for large or multimodal language models, such as Fu et al. (2025); You et al. (2024), emphasize factual accuracy or instruction following, seldom addressing symbolic interpretation or epistemic alignment. Recent cultural evaluation efforts like M3Exam Zhang et al. (2023) and SEED-Bench Li et al. (2024), although primarily focused on multimodal understanding with limited cultural coverage, begin to incorporate cultural knowledge but focus on factual recall rather than interpretive reasoning. ArtGPT Yuan et al. (2024), for instance, evaluates stylistic generation but lacks formal metrics for interpretive depth. While prior work explores aesthetic reasoning Wang (2024), these studies rarely offer structured, multi-capability evaluation. Our work addresses this gap by introducing cultural adaptability, operationalized through a multi-dimensional human expert benchmark with capability rubrics, enabling quantitative comparison in high-context domains like Chinese art.
@@
Gap in Multimodal Input–Interpretation Evaluation.
@@
Surveys Guo et al. (2023) discuss LLM limitations in nuanced discourse, but few studies compare model responsiveness to structured cultural interventions. Our empirical evaluation shows persona and knowledge base intervention improves symbolic reasoning and argumentative coherence by over 20–30%, highlighting epistemic alignment’s role beyond fluency. This cross-model, capability-specific analysis distinguishes our work.
@@ 3.1 MHEB Construction and Annotation Process
(1) Chinese Art Theory: Building on Xie He’s Six Canons (六法, 550 CE) Xie (550), particularly the concepts of “spirit resonance” (气韵生动) and “bone method” (骨法用笔), which inform our Feature Focus dimension’s emphasis on brushwork technique, artistic conception, and emotional expression.
@@
Data Collection. We collected 163 expert commentaries from authoritative sources including museum catalogs from the Palace Museum Beijing, National Palace Museum Taipei, and Metropolitan Museum of Art, as well as peer-reviewed art history journals and monographs by recognized scholars specializing in Qing court painting. Each commentary averages 500-800 Chinese characters and provides in-depth analysis of specific paintings from the “Twelve Months” series. The annotation process generated 558 total annotation instances (163 texts × 3 annotators plus quality control samples), which were consolidated into 163 final records after resolving disagreements.
Expert Sources. The 163 commentaries in MHEB were extracted from scholarly publications by 9 distinguished art historians specializing in Chinese painting and Qing court art. The corpus includes: Xue Yongnian (薛永年, 17 texts from two monographs), Wang Di (汪涤, 28 texts), Yang Danxia (杨丹霞, 28 texts), Nie Chongzheng (聂崇正, 15 texts), Shan Guoqiang (单国强, 18 texts), Li Shi (李湜, 17 texts), Xu Jianrong (徐建融, 17 texts), Zhu Wanzhang (朱万章, 11 texts), and Chen Yunru (陈韵如, 12 texts). These experts represent major institutions including the Palace Museum Beijing, National Palace Museum Taipei, and leading Chinese art history departments, ensuring diverse yet authoritative perspectives on Giuseppe Castiglione’s “Twelve Months” series.
@@
Inter-Annotator Agreement (IAA). To quantitatively assess the reliability of our annotations, we calculated inter-annotator agreement using two complementary metrics. For categorical labels (e.g., stance categories), we computed Fleiss’ kappa Fleiss (1971), which measures agreement beyond chance for multiple raters. For continuous scores (e.g., feature prominence ratings from 0-1), we calculated the intraclass correlation coefficient (ICC) Shrout and Fleiss (1979), which assesses the consistency of quantitative measurements across raters. The average Fleiss’ kappa across stance categories was 0.78, indicating substantial agreement according to Landis and Koch’s interpretation scale. The ICC for feature prominence scores reached 0.82, demonstrating excellent reliability. When disagreements occurred (defined as $\kappa$ << 0.6 for specific labels), they were resolved through discussion, with a senior art historian serving as arbiter for persistent conflicts. The stable inter-rater agreement ($\kappa$ variation << 0.05 across time) validated the effectiveness of our quality control measures. Final dataset statistics show balanced representation across different evaluative stances (Historical: 31%, Aesthetic: 28%, Technical: 23%, Comparative: 18%) and comprehensive coverage of feature focus.
@@ 3.2 Feature Engineering from Human Expert Critiques
Framework foundation relies on human expert commentaries, significantly from Giuseppe Castiglione’s (Lang Shining, 郎世宁) “Twelve Months” (十二月令图) series—Qing imperial court paintings fusing Chinese and Western traditions. To enhance model training and evaluation, a sliding window cropping strategy (640×640 pixel sub-images) was applied to these high-resolution images, augmenting data diversity and granularity for improved VLM detail recognition and evaluation accuracy, a common practice in computer vision (e.g., Lin et al. (2014); Krishna et al. (2017)).
@@ 3.3 Evaluation of Dimensions and Label System
The three dimensions of our framework (i.e., Evaluative Stance, Feature Focus, and Commentary Quality) were derived from multiple sources: (1) traditional Chinese painting theory, particularly Xie He’s “Six Principles of Painting” (谢赫六法) Acker (1954) which emphasizes spirit resonance (气韵生动), bone method (骨法用笔), and correspondence to nature (应物象形); (2) Western art criticism frameworks from Panofsky’s three levels of meaning Panofsky (1955) and Wölfflin’s formal analysis principles Wölfflin (1950); (3) consultations with curators from the Palace Museum and Metropolitan Museum who validated the relevance of these dimensions for Qing court painting analysis; and (4) empirical analysis of recurring patterns in our collected expert commentaries.
Evaluative Stance characterizes the rhetorical or evaluative position taken by the commentator (e.g., historical interpretation, praise, or critique). Feature Focus identifies the specific visual or contextual aspects discussed in the commentary (e.g., line quality, symbolism, spatial composition). Commentary Quality captures the analytical depth and logical structure of the commentary, ranging from clear, well-argued insights to superficial or biased remarks. Furthermore, each dimension comprises a set of fine-grained subcategories with bilingual English–Chinese mappings. Full definitions and label lists are provided in Appendix C.5.
@@ 3.4 Construction and Definition of Critic Personas
To capture holistic critique style and depth beyond granular features, we constructed “critic personas” representing archetypal critical perspectives. Their development was data-driven, analyzing features from human expert commentaries, complemented by art history domain expertise. Five core personas were defined: Comprehensive Analyst (博学通论型), Historically Focused Critic (历史考据型), Technique & Style Focused Critic (技艺风格型), Theory & Comparison Focused Critic (理论比较型), and General Descriptive Profile (泛化描述型). These five core personas represent data-driven evaluation categories derived from clustering analysis of human expert features, serving as benchmarks for assessing whether VLM outputs align with recognizable expert critique patterns.
@@ 3.5 Task Definition
Inputs were multifaceted: (i) high-definition “Monthly Images” (sometimes segmented); (ii) predefined “Persona Cards” (adapting the persona simulation approach from Jiang et al. (2024) to cultural domain expertise) serving as experimental interventions-distinct from the five evaluaction personas above, these eight cultural perspective prompts guided VLM generation analysis: Mama Zola (佐拉妈妈), Professor Elena Petrova (埃琳娜·佩特洛娃教授), Okakura Kakuzo (冈仓天心), Brother Thomas (托马斯修士), John Ruskin (约翰·罗斯金), Su Shi (苏轼), Guo Xi (郭熙), and Dr. Aris Thorne (阿里斯·索恩博士); (iii) standardized prompt templates Nayak et al. (2024); and (iv) optional domain knowledge resources Zhang et al. (2025b); Bin et al. (2024). Persona guidance aimed to assess VLM capability to simulate diverse perspectives and analytical styles Zhang et al. (2025a). See Appendix B for a detailed summary of each critic persona included in our study. To avoid confusion, we distinguish between the use of personas at two different levels: the five core personas described in the previous sub-section are data-driven evaluation categories for classifying generated critiques based on feature patterns, while the eight persona cards are cultural perspective prompts used to guide VLM generation during experiments. The former evaluates outputs, while the latter shapes inputs.
@@ 4.3 Effectiveness of Persona-Guided Interventions
1 Qwen-2.5-VL-7B + Mama Zola (佐拉妈妈) + KB 9.2/10 100%
2 meta-llama_Llama-4-Scout-17B-16E-Instruct + John Ruskin (约翰·罗斯金) + KB 8.9/10 97%
3 meta-llama_Llama-4-Scout-17B-16E-Instruct + Mama Zola (佐拉妈妈) + KB 8.7/10 95%
4 meta-llama_Llama-4-Scout-17B-16E-Instruct + Brother Thomas (托马斯修士) + KB 8.5/10 92%
5 meta-llama_Llama-4-Scout-17B-16E-Instruct + Su Shi (苏轼) + KB 8.5/10 92%
@@ A.1 Lang Shining’s "Twelve Months" Dataset
Our study centers on Giuseppe Castiglione’s “Twelve Months” series (十二月令图), 12 paintings showing seasonal activities in the Qing imperial court. These paintings fuse Chinese and Western artistic traditions, ideal for cross-cultural interpretation study. We compiled digital images (6 million pixels) from the National Palace Museum (Taiwan) digital archives under CC BY 4.0 license. The dataset includes historical texts and scholarly analyses in both Chinese and English, from Qing Dynasty sources and modern scholarship.
@@ Appendix B Persona Definitions
B.1 Mama Zola (佐拉妈妈)
@@
B.2 Okakura Kakuzo (冈仓天心)
@@
B.3 Professor Elena Petrova (埃琳娜·佩特洛娃教授)
@@
B.4 Brother Thomas (托马斯修士)
@@
B.5 John Ruskin (约翰·罗斯金)
@@
    • “To see clearly is poetry, prophecy, and religion, — all in one.”
B.6 Su Shi (苏轼)
@@
  • Key Influences/Background: Deeply influenced by Confucianism, Daoism, and Chan (Zen) Buddhism. Advocated for “scholar-official painting” (士人画), emphasizing the integration of poetry, calligraphy, and painting, and the expression of inner spirit. His artistic ideas had a profound impact on the development of later literati painting.
  • Analytical Style and Characteristics: Values the “spiritual resonance” (神韵) and “artistic interest” (意趣) of artworks over external formal likeness. Emphasizes the decisive role of the artist’s personal character, knowledge, and cultivation in creation. Esteems an aesthetic realm of natural innocence, plainness, and distanced simplicity.
@@
B.7 Guo Xi (郭熙)
  • Basic Information: Outstanding Chinese Northern Song Dynasty landscape painter and painting theorist (male, c. 1023-c. 1085, Wen County, Heyang). Served as an Erudite (艺学) in the imperial painting academy during Emperor Shenzong’s reign.
  • Key Influences/Background: Inherited and developed the traditions of the Northern school of landscape painting, emphasizing observation and experience of nature. His theoretical work “The Lofty Message of Forests and Streams” (林泉高致) is a seminal text in Chinese landscape painting theory.
  • Analytical Style and Characteristics: Emphasized that landscape paintings should be “walkable, viewable, wanderable, and habitable” (可行、可望、可游、可居). Proposed methods for observing and depicting landscapes such as the “Three Distances” (三远): high distance (高远), deep distance (深远), level distance (平远). Valued the influence of seasons and climate on scenery, striving for majestic and varied artistic conceptions (意境).